Calculate the molality of a 19% HF aqueous solution. What is the ppm of HF in this solution?


Determine the mass percentage and ppm of a 0.333m NaCl solution.


Calculate the molarity of a 0.85m lithium chloride aqueous solution having a density of 1.44 g/mL


1
Expert's answer
2021-05-27T07:33:59-0400

Calculate Molality of HF

1) Let us assume 100.0 gram of solution. Therefore:

19gm is HF

81gm is H2O


2) Molality is:

moles HF = 19gm/20.0059 g/mol = 0.95 mol

kg of water = 0.0810 kg

"molality = \dfrac{moles\ solute}{kilograms\ solvent}"


mololity = 0.95 mol/0.0810 kg = 1.172 molal


ppm="\dfrac{19}{1000}\times 1000000"

= 19000ppm


Calculate Mass Percentage of NaCl

1) moles of NaCl Solution given = 0.333 mol

molecular weight of NaCl = 0.333 "\times" 58.44

= 19.46 g

2) so mass percentage of NaCl is "= \dfrac{molecular\ weight}{Total\ solution}"

so "= \dfrac{19.46}{1000}\times 100"

mass percentage= 1.946% NaCl.

ppm = "\dfrac{19.46}{1000}\times 1000000"


= 19460ppm

Calculate Molarity of lithium chloride

1) Molarity in terms of molality

"molarity=\dfrac{molality\times(density\times1000 - molar\ mass)}{1000}"

so, "molarity=\dfrac{0.85\times(1.44\times1000 - \ 42.394)}{1000}"

solve above equation

molarity = 1.187 M